### v1

# Integrated GARCH-GRU in Financial Volatility Forecasting

Abstract
In this study, we propose a novel integrated Generalized Autoregressive Conditional Heteroskedasticity–Gated Recurrent Unit (GARCH-GRU) model for financial volatility modeling and forecasting. The model embeds the GARCH(1,1) formulation directly into the GRU cell architecture, yielding a unified recurrent unit that jointly captures both traditional econometric properties and complex temporal dynamics. This hybrid structure leverages the strengths of GARCH in modeling key ‘stylized facts’ of financial volatility, such as clustering and persistence, while utilizing the GRU’s capacity to learn nonlinear dependencies from sequential data. Compared to the GARCH-LSTM counterpart, the GARCH-GRU model demonstrates superior computational efficiency, requiring significantly less training time, while maintaining and improving forecasting accuracy. Empirical evaluation across multiple financial datasets confirms the model’s robust outperformance in terms of mean squared error (MSE) and mean absolute error (MAE) relative to a range of benchmarks, including standard neural networks, alternative hybrid architectures, and classical GARCH-type models. As an application, we compute Value-at-Risk (VaR) using the model’s volatility forecasts and observe lower violation ratios, further validating the predictive reliability of the proposed framework in practical risk management settings.
Keywords: Volatility forecasting, Deep Learning, Neural Networks, GARCH models, Gated Recurrent Units, Long Short Term Memory, Value at Risk, Risk Management.
1 Introduction
Volatility modeling and forecasting have long stood at the forefront of financial econometrics, underpinning crucial tasks such as risk management, derivatives pricing, portfolio optimization, and regulatory compliance. As a fundamental measure of financial market uncertainty and risk, volatility plays a central role in both academic inquiry and industry applications. Accurate volatility prediction is indispensable for understanding market dynamics, pricing assets under uncertainty, and managing financial exposure.
Financial time series exhibit several empirically observed patterns, often referred to as ‘stylized facts’ of volatility. These include volatility clustering, long memory, asymmetric responses to shocks (leverage effects), heavy-tailed return distributions, and mean-reverting behavior. Capturing these features is essential for any model aiming to provide realistic and reliable forecasts. Among these, volatility clustering and persistence are particularly well-documented: periods of high volatility tend to be followed by high volatility, and low-volatility periods likewise tend to persist. These temporal dependencies contradict the assumptions of constant variance and i.i.d. returns commonly assumed in classical models, motivating the development of more sophisticated volatility modeling approaches. Another key empirical feature is asymmetry, where negative price shocks tend to have a larger impact on future volatility than positive ones of similar magnitude. This leverage effect further challenges symmetric and linear modeling assumptions.
Traditional econometric approaches, particularly the Autoregressive Conditional Heteroskedasticity (ARCH) model introduced by Engle (1982) and its generalization, the Generalized ARCH (GARCH) model proposed by Bollerslev (1986), laid the foundation for modern volatility modeling. These models, grounded in strong statistical theory, have proven effective in modeling time-varying conditional variance and capturing volatility clustering. Their interpretability, tractability, and widespread applicability have established them as benchmarks in financial econometrics. However, GARCH-family models rely on linear parametric structures and are often limited in their ability to accommodate non-linear dynamics, structural breaks, and extreme events commonly observed in financial data (Hansen and Lunde (2005)). Several extensions, including asymmetric GARCH variants, long-memory models such as FIGARCH, and more recently, fractional affine multi-component GARCH frameworks, have sought to address some of these limitations. For example, Augustyniak et al. (2022) integrate the Heston-Nandi option pricing framework with FIGARCH dynamics to jointly capture both short- and long-memory components in volatility evolution.
With the advent of high-frequency data and advances in computational techniques, machine learning and deep learning models have emerged as powerful alternatives for volatility forecasting. Models such as recurrent neural networks (RNNs), long short-term memory (LSTM) networks (Hochreiter and Schmidhuber (1997)), and gated recurrent units (GRUs) (Chung et al. (2014)) offer a data-driven approach to uncover complex, nonlinear temporal dependencies in financial time series. These neural architectures have demonstrated superior forecasting accuracy in many settings by flexibly adapting to changing market regimes and capturing intricate dependencies that traditional econometric models may miss. Despite their empirical success, however, deep learning models often suffer from limited interpretability and lack the theoretical grounding inherent in econometric models. Moreover, the use of LSTM-based volatility models in practice is often hampered by computational inefficiency and over-parameterization, especially in latency-sensitive applications such as high-frequency trading.
Recent efforts have attempted to bridge the gap between statistical rigor and data-driven flexibility by combining GARCH-type models with deep learning architectures. Most existing hybrid models have focused on integrating GARCH with LSTM networks. While such hybrids have demonstrated improved predictive performance, the substantial computational overhead associated with LSTMs, which stem from their intricate gating mechanisms and high parameter count, poses practical limitations, particularly in latency-sensitive applications such as high-frequency trading. In contrast, the GRU model offers a more computationally efficient alternative due to its streamlined architecture, while retaining strong capabilities in modeling temporal dependencies (Chung et al. (2014)). Moreover, many existing hybrid approaches adopt a static architecture in which GARCH-generated residuals or conditional variances are simply fed as exogenous inputs to neural networks. Such designs fail to fully exploit the potential for dynamic and reciprocal interactions between econometric structures and learned representations. A notable advancement by Zhao et al. (2024) proposes embedding GARCH regression directly into LSTM cells, effectively replacing the output gate and allowing for deeper integration of volatility dynamics. However, the use of scalar hidden states and the implicit assumption of zero-mean innovations in their model limit its ability to capture external market influences and the complex, high-dimensional temporal dependencies characteristic of modern financial environments.
This study addresses the aforementioned research gaps by introducing a novel integrated GARCH-GRU framework that unifies the interpretability and statistical rigor of traditional econometric models with the nonlinear learning capabilities of deep neural networks. Specifically, we design an architecture that embeds GARCH(1,1) dynamics directly into GRU cells, preserving the gating mechanisms and hidden state transitions intrinsic to the GRU structure. This embedded design enables the model to jointly learn key volatility ‘stylized facts’, such as clustering and persistence, alongside complex temporal patterns in an end-to-end training paradigm, where both the GARCH parameters and GRU weights are optimized simultaneously. The architecture further supports hierarchical feature fusion through stacked integrated layers, facilitating multi-scale representation learning that captures heterogeneous dynamics across time horizons. To evaluate the robustness and generalizability of the proposed framework, we conduct extensive empirical experiments on a diverse set of financial assets, including the S&P 500 Index, the Dow Jones Industrial Average (DJI), and the NASDAQ Composite Index, under varying market regimes and volatility conditions.
The rest of this paper is structured as follows. Section 2 provides a detailed review of related literature in volatility modeling, spanning both econometric and machine learning paradigms. Section 3 presents the methodological framework, including the preliminaries and theoretical formulation of the GARCH-GRU model. Section 4 details the empirical study, describing the datasets, risk metrics, experimental setup, and comparative evaluation against benchmark models. Section 5 concludes the paper by summarizing key findings, discussing limitations, and outlining future directions for extending the proposed framework.
2 Related Works
Classical econometric models, such as the Autoregressive Conditional Heteroskedasticity (ARCH) model introduced by Engle (1982) and its generalization, the Generalized Autoregressive Conditional Heteroskedasticity (GARCH) model by Bollerslev (1986), laid the groundwork for capturing time-varying volatility and volatility clustering in asset returns. These models are grounded in a well-defined probabilistic framework and have served as the benchmark in both academic and applied settings. To accommodate empirical regularities beyond volatility clustering, numerous extensions to the basic GARCH framework have been introduced. The Integrated GARCH (IGARCH) model by Engle and Bollerslev (1986) incorporates a unit root in the conditional variance process, allowing for shocks to have persistent effects on volatility. The GARCH-in-Mean (GARCH-M) model proposed by Engle et al. (1987) explicitly incorporates conditional variance into the mean equation, capturing the risk-return trade-off. To address the observed asymmetry in volatility responses, where negative returns often lead to larger increases in volatility than positive ones, Nelson (1991) introduced the Exponential GARCH (EGARCH) model, which models the logarithm of the variance and naturally accommodates asymmetric effects. Similarly, the GJR-GARCH model by Glosten et al. (1993) includes an indicator function to differentiate the impact of positive and negative shocks, while the Threshold GARCH (TGARCH) model developed by Zakoian (1994) models the conditional standard deviation and also captures leverage effects. Recognizing the limitations of univariate models in capturing interdependencies across multiple financial time series, multivariate GARCH-type models have also been developed. For instance, the bivariate ARCH model introduced by Engle et al. (1984) represents an early attempt to model co-movements in volatility. The Baba-Engle-Kraft-Kroner (BEKK) model formulated by Engle and Kroner (1995) provides a flexible yet computationally tractable framework for multivariate volatility dynamics. Further extensions include the Varying Correlation MGARCH (VC-MGARCH) model proposed by Tse and Tsui (2002), which allows for time-varying conditional correlations, and the regime-switching MGARCH specification of So and Tse (2009), which accounts for nonlinear shifts in volatility regimes and better accommodates extreme events and tail risk behavior in multivariate settings.
While these models offer a robust and interpretable framework, their inherent reliance on linear dynamics and predefined functional forms limits their capacity to capture complex nonlinear relationships, structural breaks, and regime shifts commonly observed in financial markets. Moreover, their parametric nature poses challenges in adapting to rapidly evolving market conditions, particularly under high-frequency and large-scale data environments.
In response to these limitations, deep learning approaches have been increasingly applied to volatility modeling and forecasting. Recurrent neural networks (RNNs), especially Long Short-Term Memory (LSTM) networks and Gated Recurrent Units (GRUs), have demonstrated strong capabilities in learning long-range dependencies and nonlinear dynamics in time series data. The LSTM model, introduced by Hochreiter and Schmidhuber (1997) and further extended by Gers et al. (2000) and Fischer and Krauss (2018), has shown empirical superiority in capturing sequential patterns in financial datasets. In a comparative study, Liu (2019) found that LSTM-based models outperform traditional GARCH variants for volatility prediction over longer forecast horizons. GRU, empirically validated by Chung et al. (2014), offers a more computationally efficient alternative to LSTM while maintaining comparable performance in sequence modeling tasks. It has been successfully employed in various financial prediction tasks, including the Two-Stream GRU (TGRU) model proposed by Lien Minh et al. (2018), which integrates bidirectional recurrent mechanisms and achieves improved predictive accuracy over classical models.
Despite their predictive power, deep neural networks often lack financial interpretability and theoretical grounding. Their ‘black-box’ nature poses challenges for transparency, risk attribution, and regulatory acceptability. Furthermore, they do not explicitly model conditional heteroskedasticity, and thus may overlook important volatility-specific features that are well captured by econometric models.
To reconcile the statistical interpretability of GARCH with the representational flexibility of deep learning, a growing body of literature has explored hybrid approaches. Early efforts typically relied on sequential or decoupled modeling pipelines, where GARCH models were used to extract residuals or conditional variances, which were then passed as inputs into neural networks for downstream prediction. Notable examples of such decoupled architectures include the hybrid models developed by Roh (2007), Tseng et al. (2008), Hajizadeh et al. (2012), and Kristjanpoller et al. (2014); Kristjanpoller and Minutolo (2016); Kristjanpoller and Hernández (2017), as well as Kim and Won (2018). While effective in improving forecasting accuracy, these approaches typically involve separate training processes for the statistical and neural components, limiting their capacity to learn joint representations and often resulting in inefficiencies and suboptimal integration. Recent studies have proposed end-to-end hybrid architectures that attempt to jointly optimize econometric and deep learning components. For instance, Zhao et al. (2024) introduced an embedded GARCH-LSTM model that integrates the GARCH regression mechanism within LSTM cells by replacing the output gate. This approach represents a significant step toward unified modeling, enabling direct interaction between GARCH dynamics and temporal feature extraction. However, this model imposes strong assumptions, such as scalar hidden state representations and zero-mean innovations, which restrict its expressiveness and ability to model complex market behaviors and external covariate influences.
While substantial progress has been made in both traditional volatility modeling and machine learning-based forecasting, existing approaches either lack the interpretability required for financial applications or fail to fully utilize the learning capacity of neural networks. The current literature also tends to emphasize LSTM-based hybrid models, leaving the advantages of GRU-based integrations largely untapped. This motivates the development of new unified frameworks that can simultaneously capture volatility stylized facts and complex temporal dependencies in a computationally efficient and theoretically coherent manner.
3 Methodology
3.1 Theoretical Foundations
3.1.1 GARCH Model
The GARCH model (Bollerslev (1986)) includes lagged conditional variances along with lagged squared residuals, effectively capturing volatility clustering and persistence in financial time series. Assuming a constant mean return and given the information set $\psi_{t-1}$, the model is specified as:
$r_{t}=\mu+\epsilon_{t}$ (1)
$\epsilon_{t}|\psi_{t-1}\sim\mathcal{D}(0,\sigma_{t}^{2})$ (2)
$\sigma_{t}^{2}=\omega_{0}+\sum_{i=1}^{p}\alpha_{i}\epsilon_{t-i}^{2}+\sum_{j=1}^{q}\beta_{j}\sigma^{2}_{t-j}$ (3)
where $\mathcal{D}(0,\sigma_{t}^{2})$ typically denotes a normal or Student’s t-distribution. The model parameters satisfy $\omega>0$, $\alpha_{i}\geq 0$, $\beta_{j}\geq 0$, and $\sum_{i=1}^{p}\alpha+\sum_{j=1}^{q}\beta<1$ to ensure positivity and weak stationarity of the conditional variance process. The coefficients $\alpha_{i}$ capture the short-term impact of shocks, while the coefficients $\beta_{j}$ reflect the persistence of volatility over time.
3.1.2 Recurrent Neural Networks
Both LSTM and GRU models are widely adopted recurrent neural network (RNN) architectures for sequential modeling and have demonstrated strong empirical performance in financial forecasting tasks. These models are particularly effective at capturing temporal dependencies in time series data due to their gating mechanisms, which regulate the flow of information through the network over time. LSTM is designed to handle long-range dependencies by introducing memory cells and three gating components: input, forget, and output gates. These gates enable the model to selectively retain or discard information from previous time steps, thereby mitigating the vanishing gradient problem associated with standard RNNs. The LSTM cell computes the hidden state and cell state at time $t$ through the following set of equations:
$f_{t}$ $=\sigma\left(W_{f}x_{t}+U_{f}h_{t-1}+b_{f}\right)$ (4)
$i_{t}$ $=\sigma\left(W_{i}x_{t}+U_{i}h_{t-1}+b_{i}\right)$ (5)
$o_{t}$ $=\sigma\left(W_{o}x_{t}+U_{o}h_{t-1}+b_{o}\right)$ (6)
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}_{t}$ (7)
$\tilde{c}_{t}$ $=\tanh\left(W_{c}x_{t}+U_{c}h_{t-1}+b_{c}\right)$ (8)
$h_{t}$ $=o_{t}\odot\tanh\left(c_{t}\right)$ (9)
where $x_{t}$ is the input vector at time $t$, $h_{t-1}$ is the previous hidden state, $c_{t-1}$ is the previous cell state, and $\sigma(\cdot)$ denotes the sigmoid activation function. The memory cell $c_{t}$ acts as an internal state that accumulates information over time, while $h_{t}$ serves as the output at time $t$. The element-wise product is denoted by $\odot$.
The GRU is a streamlined variant of the LSTM that retains the key advantages of gated memory while reducing computational complexity. It employs only two gates, the update gate and reset gate, thereby simplifying the architecture and accelerating training. The GRU computes its hidden state via the following equations:
$z_{t}$ $=\sigma\left(W_{z}x_{t}+U_{z}h_{t-1}+b_{z}\right)$ (10)
$r_{t}$ $=\sigma\left(W_{r}x_{t}+U_{r}h_{t-1}+b_{r}\right)$ (11)
$\tilde{h}_{t}$ $=\tanh\left(W_{h}x_{t}+U_{h}(r_{t}\odot h_{t-1})+b_{h}\right)$ (12)
$h_{t}$ $=\left(1-z_{t}\right)\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (13)
where $z_{t}$ and $r_{t}$ control how much past information is carried forward and how much is reset at each time step, respectively. Unlike LSTM, GRU does not maintain a separate cell state, but instead uses the hidden state $h_{t}$ for both memory and output.
3.2 Integrated GARCH-GRU Framework
3.2.1 GARCH-GRU Model
The proposed integrated GARCH-GRU model builds upon the conventional GARCH(1,1) specification with constant mean assumption in Eq.1, which is formulated as:
$\sigma_{t}^{2}=\omega_{0}+\alpha\epsilon_{t-1}^{2}+\beta\sigma_{t-1}^{2}.$ (14)
Asset returns $r_{t}$ exhibit innovations $\epsilon_{t}$ with time-varying conditional variance $\sigma_{t}^{2}$. The key innovation lies in structural integration of GARCH dynamics into GRU operations through a multi-stage gating mechanism, as formalized in Eqs.(15)-(20).
$g_{t}$ $=\phi\left(\omega_{0}+\alpha\epsilon_{t-1}^{2}+\beta\sigma_{t-1}^{2}\right)$ (15)
$z_{t}$ $=\sigma\left(W_{z}x_{t}+U_{z}h_{t-1}+b_{z}\right)$ (16)
$r_{t}$ $=\sigma\left(W_{r}x_{t}+U_{r}h_{t-1}+b_{r}\right)$ (17)
$\tilde{h}_{t}$ $=\tanh\left(W_{h}x_{t}+U_{h}\left(r_{t}\odot h_{t-1}\right)+b_{h}\right)$ (18)
$\hat{h}_{t}$ $=\left(1-z_{t}\right)\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}$ (19)
$h_{t}$ $=\tanh\left(\hat{h}_{t}+\gamma g_{t}\right)$ (20)
The GARCH-GRU cell operates through dual processing streams. First, the GARCH component generates volatility estimates $g_{t}$ using recombination of past residuals and variances (Eq.15), where $\phi(\cdot)=W_{g}g_{t}+b_{g}$ denotes a linear projection aligning scalar GARCH output with GRU hidden state dimensions. Second, a modified GRU architecture processes sequential inputs through standard reset/update gates (Eqs.(16)-(19)), followed by additive integration of GARCH signals into the hidden state update (Eq.(20)). The learnable parameter $\gamma$ governs the contribution intensity of GARCH components, while element-wise addition enables granular modulation of individual hidden units. This design preserves GRU’s capacity to model nonlinear temporal dependencies while explicitly encoding volatility persistence patterns. A graphical representation of the integrated GARCH-GRU framework is shown in Figure.1.
Figure 1: Graphical Representation of integrated GARCH-GRU
To ensure stability of the GARCH component, we apply a reparameterization to enforce the stationarity constraints on parameters ($\omega_{0}$, $\alpha$, $\beta$), especially for $\alpha+\beta<1$. The GRU subsystem employs dropout regularization to prevent overfitting while maintaining gradient flow through deep layers. Final volatility predictions are generated through a constrained output layer that applies softplus activation followed by square-root transformation, ensuring non-negative volatility estimates consistent with financial time series properties.
This integrated framework achieves three complementary objectives: explicit representation of volatility clustering through GARCH components, adaptive extraction of complex temporal dependencies via GRU networks, and preservation of econometric validity through interpretable parameter constraints. The additive integration mechanism formalized in Eq.(20) establishes a mathematically grounded interface between autoregressive variance dynamics and learned neural representations. This synergistic combination enables joint optimization of statistical and deep learning components while retaining structural interpretability, ensuring that volatility forecasts simultaneously capture conventional persistence patterns and nonlinear feature interactions inherent in financial time series.
3.2.2 GARCH-LSTM Model
To facilitate a systematic comparison between recurrent architectures in hybrid volatility modeling, we construct an integrated GARCH-LSTM model that parallels the design of the proposed GARCH-GRU framework, as formalized in Eqs.(21)–(28). Consistent with the GARCH-GRU’s integration strategy, the GARCH-LSTM model introduces a volatility-driven multiplicative gate derived from the standard GARCH(1,1) specification (Eq.(21)), which encodes key stylized facts of volatility, including clustering and persistence.
While both architectures incorporate GARCH signals to modulate neural dynamics, they diverge in the final fusion mechanism. In the GARCH-GRU model, the GARCH component is additively combined with the intermediate hidden state (Eq.(20)), allowing for element-wise modulation controlled by a learnable scaling parameter. In contrast, the GARCH-LSTM model adopts a multiplicative modulation strategy, partially inspired by the integration scheme proposed in Zhao et al. (2024). Specifically, the final hidden state $h_{t}$ is obtained by scaling the LSTM’s intermediate output $\tilde{h}_{t}$ with a learnable coupling parameter $w$ and a nonlinear transformation of the GARCH component $g_{t}$ (Eq.(28)). This multiplicative gating mechanism allows $w$ to adaptively calibrates the magnitude of GARCH contributions relative to the LSTM’s intrinsic temporal dynamics. As $w$ increases, the magnitude of $g_{t}$ becomes larger, thereby intensifying its modulation effect on the final hidden state and increasing its contribution to the model’s volatility forecast.
$g_{t}$ $=\omega_{0}+\alpha\epsilon_{t-1}^{2}+\beta\sigma_{t-1}^{2}$ (21)
$f_{t}$ $=\sigma\left(W_{f}x_{t}+U_{f}h_{t-1}+b_{f}\right)$ (22)
$i_{t}$ $=\sigma\left(W_{i}x_{t}+U_{i}h_{t-1}+b_{i}\right)$ (23)
$o_{t}$ $=\sigma\left(W_{o}x_{t}+U_{o}h_{t-1}+b_{o}\right)$ (24)
$\tilde{c}_{t}$ $=\tanh\left(W_{c}x_{t}+U_{c}h_{t-1}+b_{c}\right)$ (25)
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot\tilde{c}_{t}$ (26)
$\tilde{h}_{t}$ $=o_{t}\cdot tanh(c_{t})$ (27)
$h_{t}$ $=\left(1+w\cdot\tanh\left(g_{t}\right)\right)\odot\tilde{h}_{t}$ (28)
Unlike the GARCH-GRU’s element-wise additive fusion, which locally adjusts individual hidden units, the GARCH-LSTM’s approach globally amplifies or attenuates the entire hidden state proportionally to the volatility estimate $g_{t}$. And the $tanh(\cdot)$ normalization ensures bounded scaling while preserving the directional consistency of hidden state updates, thereby enabling the model to emphasize high-volatility regimes without destabilizing gradient propagation.
4 Empirical Study
4.1 Experimental Design
Datasets: We collect daily closing prices $p_{t}$ of three major U.S. stock indices, S&P 500, Dow Jones Industrial Average (DJI), and NASDAQ Composite, over the period from January 1, 2010, to December 31, 2019. This decade-long sample spans various market conditions, including periods of both low and high volatility. To construct the return series, we follow standard practices in the financial econometrics literature (Roh (2007), Bucci (2017), Kim and Won (2018), Bucci (2020)) and compute daily log returns as:
$r_{t}=log{\frac{p_{t}}{p_{t-1}}}$ (29)
The realized volatility used as the target variable for forecasting is computed using a rolling window of $k=5$ days (approximately one trading week) as follows:
$\sigma_{t}$ $=\sqrt{\frac{1}{k}\sum_{i=0}^{k}(r_{t-i}-\mu_{k})^{2}}$ (30)
$\mu_{k}$ $=\frac{1}{k}\sum_{i=0}^{k}r_{t-i}$
The last 252 trading days, corresponding to roughly one calendar year, are used as the test set. Within the remaining training set, the final 20% is reserved for validation during model training and tuning.
Table 1: Statistics for Dataset
Dataset Count Mean Standard deviation Skewness Kurtosis ADF Non-zero Mean t-test LM
S&P 500 2414 0.0406 0.9307 -0.5746 4.5604 $-24.5061^{**}$ $2.1439^{*}$ $432.4055^{**}$
DJI 2414 0.0397 0.8858 -0.5697 3.9235 $-24.2261^{**}$ $2.2038^{*}$ $405.4800^{**}$
NASDAQ 2414 0.0501 1.0733 -0.5236 3.3275 $-24.0755^{**}$ $2.2947^{*}$ $342.0940^{**}$
‘*’ indicates significance at the 5% level, and ‘**’ indicates significance at the 1% level. ‘ADF’ represents the ADF test for unit root existence, ‘LM’ represents the ARCH Lagrange Multiplier test for heteroscedasticity existence.
Table.1 presents summary statistics and diagnostic tests for the return series of each index. All return series reject the null hypothesis of the Augmented Dickey-Fuller (ADF) test at the 1% significance level, confirming that the return series are stationary and do not contain unit roots. Additionally, the Lagrange Multiplier (LM) test for ARCH effects strongly rejects the null hypothesis of homoscedasticity for all three indices, indicating the presence of conditional heteroskedasticity, and thereby justifying the application of GARCH-type models. Furthermore, a two-sided t-test for non-zero mean is performed to assess the validity of the constant mean assumption adopted in volatility models. The test results indicate statistical significance at the 5% level across all datasets, thereby supporting the validity of the constant mean assumption adopted in our models for the subsequent empirical analysis.
All models are trained and evaluated using a rolling window approach (see Fig.2) to simulate realistic forecasting conditions. A fixed-length window of 22 trading days (approximately one month) is used to train the model, which is then used to generate volatility forecasts for the next period. The forecast horizons considered in this study are 1-day, 3-day, and 7-day (1-week) ahead forecasts. After each forecast, the window is rolled forward by one day, and the process is repeated until the entire test set is covered. This strategy ensures consistency in evaluation and reflects the operational setting in real-time forecasting scenarios.
Figure 2: Rolling Window Methodology
Model Configuration: The proposed integrated GARCH-GRU architecture is implemented by introducing custom modifications to the standard GRU cell, embedding conditional heteroskedasticity dynamics in a manner consistent with the GARCH(1,1) specification. A parallel implementation is applied to the integrated GARCH-LSTM model, serving as a comparative counterpart. This hybrid design enables the models to simultaneously capture nonlinear temporal dependencies and time-varying volatility structures inherent in financial time series.
To optimize model performance, hyperparameter tuning is conducted using the Optuna framework, leveraging a Bayesian optimization strategy for efficient exploration of the parameter space. To prevent overfitting, an early stopping criterion is applied based on validation set performance, with a patience threshold of 10 epochs. The mean squared error (MSE) is used as the monitoring metric. Model training is carried out using the Adam optimizer, which offers adaptive learning rates and robust convergence properties. This training configuration ensures both stability during optimization and strong generalization across varying forecasting horizons, supporting the empirical robustness of the proposed hybrid volatility models.
Evaluation Framework: The forecasting performance of the proposed model is evaluated using two complementary statistical accuracy metrics: Mean Squared Error (MSE) and Mean Absolute Error (MAE). These metrics are computed over the out-of-sample forecast horizon to quantify the deviation of predicted volatility from realized volatility. The Mean Squared Error is defined as
$\mathrm{MSE}:=\frac{1}{n}\sum_{t=1}^{n}\left(\hat{\sigma}_{t}-\sigma_{t}\right)^{2}$
where $\hat{\sigma}_{t}$ and $\sigma_{t}$ denote the predicted and realized volatility at time $t$, respectively. MSE penalizes larger forecast errors more heavily due to the form of the quadratic loss function, making it particularly sensitive to extreme deviations. In contrast, the Mean Absolute Error is calculated as
$\mathrm{MAE}:=\frac{1}{n}\sum_{t=1}^{n}\left|\hat{\sigma}_{t}-\sigma_{t}\right|$
which provides a more robust measure of average forecast accuracy that is less influenced by outliers. Together, MSE and MAE offer a comprehensive view of both the magnitude and variability of forecast errors, facilitating comparative evaluation across benchmark models and robustness checks.
In addition to these statistical loss metrics, we employ Value-at-Risk (VaR) forecasting as a complementary evaluation tool to assess the practical utility of the volatility predictions in risk management. VaR violations and violation ratios are analyzed to quantify the accuracy of the forecasted volatility in capturing downside risk. This dual focus on statistical precision and financial relevance enables a rigorous and well-rounded evaluation of the proposed hybrid model.
Benchmarks: To rigorously evaluate the forecasting performance of the proposed integrated GARCH-GRU model, we compare it against four representative classes of benchmark models that are widely studied in the volatility forecasting literature. These benchmarks are selected to reflect a broad spectrum of modeling paradigms, encompassing classical econometric approaches, standard deep learning architectures, and both embedded and pipeline-based hybrid frameworks. The first benchmark class includes embedded hybrid models, specifically the architecture proposed in Zhao et al. (2024), which modifies the LSTM cell by replacing the output gate with a GARCH-type regression. This embedded integration enables the model to learn both volatility dynamics and sequential dependencies within a unified recurrent structure, facilitating end-to-end training. We refer to this model as ‘bl_GARCH_LSTM’, and designate it as the baseline hybrid benchmark, given its architectural similarity to our GARCH-GRU model and its innovative embedding of econometric dynamics within deep learning modules. The second benchmark class comprises pipeline hybrid models, where the output of a GARCH model, typically the estimated conditional variances, are used as input features for a recurrent neural network. This decoupled, two-stage integration strategy allows the GARCH component to capture short-term volatility clustering, while the LSTM component learns additional nonlinear and long-range dependencies. We denote this model as ‘pl_GARCH_LSTM’, representing a common and pragmatic approach in hybrid modeling that leverages the strengths of both statistical and machine learning techniques in a modular fashion. The third class consists of pure deep learning models, including standard GRU, LSTM, and Transformer architectures. These models are capable of learning rich temporal patterns and nonlinear dependencies, but do not incorporate any explicit econometric mechanisms for modeling volatility. As such, they serve as benchmarks for evaluating the added value of incorporating structured volatility dynamics into neural networks. The fourth and final class includes traditional econometric models, specifically GARCH(1,1) and GJR-GARCH(1,1,1). These models are well-established in the literature for capturing conditional heteroskedasticity and leverage effects in financial return series. Their inclusion provides a point of comparison for assessing the performance gains achieved through hybridization with neural networks. Together, these benchmark classes, comprising classical models, standard deep learning architectures, pipeline hybrids, and embedded hybrids, form a comprehensive evaluation framework. This allows us to position the proposed GARCH-GRU model within a broad methodological context, ensuring that observed performance improvements are both robust and interpretable, and reflect meaningful advances over existing approaches in the literature.
Architectural Validation: Prior to the construction of the integrated GARCH-GRU architecture, we conducted a rigorous validation of the custom neural network (NN) components to ensure their correctness and reliability. Since the hybrid model relies on manually implemented neural cells to allow GARCH integration at the architectural level, it was essential to verify that these custom components replicate the functional behavior of standard, well-tested NN implementations. This validation step was crucial to attribute any observed performance difference in the hybrid model solely to the intended GARCH-NN fusion, rather than unintended implementation artifacts. To assess implementation equivalence, we treated the custom and standard neural networks as two deterministic functions, $f_{1}(x)$ and $f_{2}(x)$, respectively. A set of randomly generated input vectors $x$ was passed through both networks, yielding output sequences $y_{1}=f_{1}(x)$ and $y_{2}=f_{2}(x)$. The output were then compared using a paired t-test, which evaluates whether the mean difference between paired observations from the two models is statistically significant. Formally, we tested the null hypothesis
$H_{0}:\mathbb{E}\left[y_{1}-y_{2}\right]=0$
against the alternative
$H_{1}:\mathbb{E}\left[y_{1}-y_{2}\right]\neq 0$
The results of the paired t-test are summarized in Table.2. For both GRU and LSTM architectures, the mean differences between the custom and standard implementations were effectively zero, and the corresponding p-values substantially exceeded conventional significance thresholds (e.g., $\alpha$ = 0.01). These results indicate that there is no statistically significant difference between the output of the custom and standard models.
Table 2: Paired t-test Comparing Custom and Standard GRU/LSTM Results
Model Mean Difference t-statistic p-value
GRU 0.000 $0.0195^{**}$ $0.9845^{**}$
LSTM -0.0003 $-0.7857^{**}$ $0.4339^{**}$
Significance tested at $\alpha=1\%$ and represented with ‘**’. Results indicate no statistically significant differences between custom and standard models.
These findings validate the correctness and reliability of the custom neural cell implementations. As such, the proposed GARCH-GRU model can be confidently attributed with structural improvements in forecasting performance, rather than artifacts introduced through computational discrepancies.
4.2 Model Evaluation
Table.3 presents the empirical forecasting results for all models, where performance metrics are reported as averages over 20 runs with different random seeds. The exception is the traditional GARCH models, which are deterministic and do not involve stochastic training procedures. The table compares the forecasting accuracy of the proposed integrated GARCH-GRU model against a comprehensive set of benchmarks, including various hybrid model variants, standalone deep learning architectures, and traditional econometric models. Here the proposed integrated GARCH-GRU model, denoted as ‘eh_GARCH_GRU’, represents an enhanced hybrid architecture that unifies GARCH dynamics with the GRU structure. This comparative framework allows for a robust assessment of the GARCH-GRU model’s effectiveness across different forecasting horizons and datasets. In addition, the in-sample estimation results for the GARCH component parameters across all relevant models are presented in Appendix.A, providing further insights into the statistical properties and internal dynamics of the hybrid architectures.
Table 3: Model Performance Comparison
Model eh_GARCH_GRU eh_GARCH_LSTM bl_GARCH_LSTM pl_GARCH_LSTM GRU LSTM Transformer GARCH(1,1) GJR-GARCH
Risk Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
S&P 500 1D 0.0202 (0.0012) 0.0976 (0.0047) 0.0256 (0.0020) 0.1073 (0.0055) 0.0616 (0.0129) 0.1940 (0.0236) 0.0224 (0.0024) 0.1006 (0.0067) 0.0217 (0.0014) 0.0989 (0.0046) 0.0230 (0.0022) 0.1026 (0.0056) 0.0234 (0.0062) 0.1040 (0.0164) 0.0795 0.2174 0.0863 0.2124
3D 0.0375 (0.0013) 0.1383 (0.0059) 0.0483 (0.0042) 0.1590 (0.0075) 0.0730 (0.0061) 0.2064 (0.0112) 0.0409 (0.0041) 0.1432 (0.0084) 0.0407 (0.0052) 0.1454 (0.0129) 0.0414 (0.0052) 0.1450 (0.0097) 0.0493 (0.0041) 0.1556 (0.0120) 0.1053 0.2529 0.1138 0.2519
1W 0.0785 (0.0035) 0.2077 (0.0100) 0.1180 (0.0162) 0.2534 (0.0166) 0.0822 (0.0121) 0.2177 (0.0148) 0.0896 (0.0269) 0.2159 (0.0156) 0.0983 (0.0109) 0.2302 (0.0173) 0.0802 (0.0048) 0.2132 (0.0085) 0.0799 (0.0025) 0.2082 (0.0075) 0.1635 0.3125 0.1961 0.3322
DJI 1D 0.0225 (0.0013) 0.1036 (0.0064) 0.0252 (0.0017) 0.1106 (0.0050) 0.0596 (0.0135) 0.1981 (0.0215) 0.0216 (0.0024 ) 0.1022 (0.0101) 0.0206 (0.0013) 0.1017 (0.0064) 0.0224 (0.0015) 0.1036 (0.0066) 0.0309 (0.0042) 0.1228 (0.0101) 0.0767 0.2101 0.0925 0.2250
3D 0.0360 (0.0016) 0.1388 (0.0045) 0.0490 (0.0041) 0.1591 (0.0058) 0.0676 (0.0046) 0.2001 (0.0084) 0.0376 (0.0024) 0.1389 (0.0040) 0.0333 (0.0009) 0.1330 (0.0013) 0.0385 (0.0041) 0.1406 (0.0080) 0.0520 (0.0086) 0.1650 (0.0105) 0.0979 0.2396 0.1202 0.2601
1W 0.0759 (0.0030) 0.2040 (0.0072) 0.1090 (0.0157) 0.2388 (0.0146) 0.7463 (0.4080) 0.5211 (0.1607) 0.0836 (0.0082) 0.2099 (0.0122) 0.0931 (0.0093) 0.2276 (0.0117) 0.0847 (0.0115) 0.2121 (0.0149) 0.0860 (0.0071) 0.2184 (0.0114) 0.1470 0.2906 0.1917 0.3257
NASDAQ 1D 0.0288 (0.0015) 0.1179 (0.0056) 0.0369 (0.0028) 0.1317 (0.0062) 0.4576 (1.5881) 0.3612 (0.5266) 0.0372 (0.0029) 0.1286 (0.0043) 0.0376 (0.0881) 0.1323 (0.0713) 0.0328 (0.0030) 0.1217 (0.0066) 0.0337 (0.0045) 0.1253 (0.0129) 0.1342 0.2761 0.1446 0.2766
3D 0.0572 (0.0017) 0.1717 (0.0053) 0.0670 (0.0055) 0.1869 (0.0091) 0.0723 (0.0045) 0.2040 (0.0084) 0.0698 (0.0041) 0.1909 (0.0076) 0.0695 (0.0056) 0.1879 (0.0095) 0.0607 (0.0066) 0.1764 (0.0112) 0.0780 (0.0090) 0.1984 (0.0119) 0.1659 0.3213 0.1835 0.3233
1W 0.1086 (0.0027) 0.2476 (0.0047) 0.1542 (0.0182) 0.2919 (0.0178) 0.1248 (0.0279) 0.2838 (0.0802) 0.1202 (0.0071) 0.2597 (0.0095) 0.1344 (0.0207) 0.2721 (0.0187) 0.1353 (0.0148) 0.2720 (0.0153) 0.1386 (0.0149) 0.2785 (0.0198) 0.2594 0.4017 0.3136 0.4169
‘eh_GARCH_GRU’ and ‘eh_GARCH_LSTM’ denote the proposed end-to-end integrated models. ‘bl_GARCH_LSTM’ refers to the embedded integration approach that modifies the output gate within the LSTM architecture. ‘pl_GARCH_LSTM’ and ‘pl_GARCH_GRU’ represent pipeline-based integration models, where GARCH and neural components are trained sequentially. Values in parentheses indicate standard deviations computed from 20 independent runs with different random seeds. As the GARCH(1,1) and GJR-GARCH models are deterministic and do not involve stochastic training procedures, standard deviations are not reported for these benchmarks.
In the overall comparison of the model forecasting errors across the three datasets, our integrated GARCH-GRU model consistently outperforms other benchmark models, achieving the lowest MSE and MAE values in most forecasting scenarios. This superior performance is especially evident in medium- and long-term forecasting horizons (3-day and 1-week) across the S&P 500 and NASDAQ indices. An exception arises in the case of the Dow Jones Industrial Average (DJI) at the 1-day and 3-day forecast horizons, where the GARCH-GRU model exhibits slightly higher forecasting errors than the baseline GRU model. This performance difference can be attributed to several factors. Notably, the DJI consists of a smaller and more stable group of blue-chip stocks, which typically exhibit lower short-term volatility and weaker clustering effects compared to the broader S&P 500 or the more volatility-prone, technology-driven NASDAQ. In such settings, the benefits of incorporating GARCH-style persistence into the model may be limited, and purely data-driven architectures like GRU may offer better adaptability to short-term fluctuations that are less driven by persistent volatility shocks. Nonetheless, the differences in errors are relatively small and statistically insignificant, and the integrated GARCH-GRU model still provides a more interpretable and principled framework.
When compared to the integrated GARCH-LSTM model, the GARCH-GRU consistently outperforms across all datasets and forecast horizons. On the S&P 500, for example, the GARCH-GRU model achieves lower errors at every horizon, with the largest gap observed at the 1-week forecast, where the MSE is reduced from 0.1180 to 0.0785, and MAE is reduced from 0.2534 to 0.2077. Similar improvements are seen on the NASDAQ and Dow Jones (DJI) indices. This performance gap reflects the efficiency of the GRU architecture, which, despite having a more compact gating structure, proves more effective when tightly integrated with GARCH dynamics. The GRU’s reduced parameter count may also contribute to better generalization in volatility forecasting tasks, especially under data constraints or high volatility regimes.
Beyond its direct comparison with the GARCH-LSTM model, the proposed GARCH-GRU architecture also outperforms alternative hybrid configurations, including the embedded GARCH-LSTM (bl_GARCH_LSTM), which replaces the LSTM output gate with a GARCH regression, and the pipeline-based GARCH-LSTM (pl_GARCH_LSTM), which sequentially combines GARCH and neural network components. These alternative designs often suffer from limited integration or strong constraints, thereby constraining the model’s ability to fully capture the complex interplay between temporal dynamics and volatility behavior. For instance, on the S&P 500 index at the 1-day horizon, the pipeline model (pl_GARCH_LSTM) yields an MSE of 0.0224, while the embedded model (bl_GARCH_LSTM) performs significantly worse with an MSE of 0.0616. In contrast, the GARCH-GRU achieves a lower MSE of 0.0202, demonstrating its superior forecasting precision. These results highlight the advantages of a tightly coupled, end-to-end architecture, where GARCH dynamics are structurally embedded within the GRU’s update equations. Such integration facilitates more expressive representation learning and enables the model to coherently internalize both nonlinear temporal dependencies and heteroskedastic variance patterns, leading to improved volatility forecasts.
The proposed GARCH-GRU model also demonstrates clear advantages over standalone neural network architectures, including standard GRU, LSTM, and Transformer models. Although these models are capable of learning nonlinear temporal patterns, they lack explicit mechanisms for modeling volatility ‘stylized facts’ such as clustering or long memory. The GARCH-GRU model bridges this gap by embedding statistical volatility dynamics into the hidden state transitions. On the Dow Jones (DJI) index at the 1-week horizon, the standalone GRU and LSTM models yield MSEs of 0.0931 and 0.0847, respectively, while the GARCH-GRU model reduces the error to 0.0759. While the Transformer model performs competitively on certain short-term horizons, its performance degrades over longer horizon, suggesting that its attention mechanism lacks the necessary structural bias for volatility modeling. These results underline the value of combining econometric priors with neural architectures to improve robustness and long-horizon predictive power.
In comparison with the traditional econometric models, such as GARCH(1,1) and GJR-GARCH, the GARCH-GRU model achieves substantial improvements across all datasets and forecasting horizons. The classical models, while interpretable and grounded in financial theory, are limited in their capacity to capture nonlinear dependencies and complex temporal structures. Across all datasets and horizons, the GARCH-GRU model achieves large reductions in both MSE and MAE. For example, on the S&P 500 dataset for 1-day forecasts, GARCH(1,1) and GJR-GARCH models report MSEs of 0.0795 and 0.0863, respectively, compared to 0.0202 from the proposed model, which reaches a reduction of over 70%. Similar margins can be seen in longer-horizon forecasts and other datasets, underscoring the limitations of purely statistical models in modern financial environments characterized by nonlinear and regime-dependent volatility dynamics.
Furthermore, to assess computational efficiency, we measured the average training time for both the integrated GARCH-GRU and GARCH-LSTM models across 20 runs with different random seeds. The results show that the GARCH-GRU model requires only 133.52 seconds on average, compared to 351.57 seconds for the GARCH-LSTM model. It indicates that the GARCH-GRU model achieves approximately 62% reduction in training time. This substantial reduction in training time highlights the computational efficiency of the GRU architecture, which benefits from a simpler gating mechanism while preserving strong sequence modeling capabilities. The ability of the GARCH-GRU model to achieve superior forecasting accuracy with significantly lower computational cost further underscores the advantage of integrating GARCH dynamics with GRU. This makes the proposed model particularly appealing for large-scale or real-time financial forecasting applications, where both performance and efficiency are critical.
5 Applications
One of the most important practical applications of volatility forecasting lies in financial risk management, particularly in the computation of Value-at-Risk (VaR). VaR is a widely adopted quantitative risk measure that estimates the maximum potential loss of a portfolio or asset over a specified time horizon at a given confidence level. It is extensively used by financial institutions, regulators, and portfolio managers to assess market risk exposure, allocate capital, and comply with regulatory requirements such as those imposed under the Basel Accords.
Formally, for a return series ${r_{t}}$, the one-period ahead VaR at significance level $\alpha\in(0,1)$ is defined as the quantile of the conditional return distribution such that the probability of a loss exceeding the VaR level is no greater than $\alpha$. Mathematically, the VaR at level $\alpha$, denoted $\text{VaR}_{t+1}^{\alpha}$, satisfies:
$\mathbb{P}(r_{t+1}<-\text{VaR}_{t+1}^{\alpha}\mid\mathcal{F}_{t})=\alpha$ (31)
where $\mathcal{F}_{t}$ represents the information set available at time $t$. Under the assumption that the return series have constant mean $\mu$, in this study, rather than assuming a parametric distribution (e.g., Gaussian), we estimate VaR using a non-parametric approach based on the empirical distribution of standardized returns. Specifically, we compute:
$\operatorname{VaR}_{t+1}^{\alpha}=\mu-q_{\alpha}(\varepsilon)\cdot\hat{\sigma}_{t+1}$ (32)
where $q_{\alpha}(\varepsilon)$ denotes the $\alpha$-quantile of the empirical distribution of $\varepsilon_{t+1}$, estimated from historical standardized residuals $\left\{\hat{\varepsilon}_{t}\right\}=\left\{\left(r_{t}-\mu\right)/\sigma_{t}\right\}$. And $\hat{\sigma}_{t+1}$ is the one-day-ahead forecasted volatility produced by the model. This method allows the VaR estimate to reflect the empirical distributional features of asset returns, such as skewness and fat tails, without imposing restrictive distributional assumptions.
To evaluate the quality of the VaR forecasts, we compute two diagnostic metrics: the violation count and the violation ratio. A violation occurs when the actual return falls below the estimated VaR threshold:
$\mathbb{I}\left\{r_{t+1}<-\operatorname{VaR}_{t+1}^{(\alpha)}\right\}=1$ (33)
indicating an underestimation of the true downside risk. The violation count is the total number of such instances, and the violation ratio is computed as:
[…]
which is compared to the nominal level $\alpha$ to assess whether the model underestimates or overestimates risk.
Figure 3: Actual Returns, VaR Forecasts, and Violations at 95% Confidence Level on S&P 500
We evaluate and compare the 1-day-ahead VaR performance of seven volatility forecasting models: the proposed integrated GARCH-GRU and GARCH-LSTM models, the embedded hybrid variant (bl_GARCH_LSTM), two pipeline-based architectures (pl_GARCH_GRU and pl_GARCH_LSTM), and two standard econometric baselines: GARCH(1,1) and GJR-GARCH. All models are evaluated on the S&P 500 dataset. The actual returns and associated VaR forecasts are illustrated in Figure 3, and the violation statistics are summarized in Table 4.
Table 4: VaR Forecast Violations (One-Day Ahead Forecasts on S&P 500)
Model Violation Numbers Violation Ratio
eh_GARCH_GRU 3 1.30%
eh_GARCH_LSTM 8 3.48%
bl_GARCH_LSTM 15 6.52%
pl_GARCH_GRU 5 2.17%
pl_GARCH_LSTM 7 3.04%
GARCH(1,1) 18 7.83%
GJR-GARCH 17 7.39%
‘eh_GARCH_GRU’ and ‘eh_GARCH_LSTM’ denote the proposed end-to-end integrated models. ‘bl_GARCH_LSTM’ refers to the embedded integration approach that modifies the output gate within the LSTM architecture. ‘pl_GARCH_LSTM’ and ‘pl_GARCH_GRU’ represent pipeline-based integration models, where GARCH and neural components are trained sequentially.
Among all models, the integrated GARCH-GRU (eh_GARCH_GRU) model achieves the lowest violation count (3) and violation ratio (1.30%). In contrast, the baseline econometric models (GARCH(1,1) and GJR-GARCH) exhibit much higher violation ratios, 7.83% and 7.39%, respectively, indicating significant underestimation of downside risk. The other hybrid models also tend to generate more violations than the GARCH-GRU, suggesting less accurate tail risk modeling. These findings underscore the practical utility of the proposed GARCH-GRU model in risk-sensitive applications, particularly in capturing complex, nonlinear volatility dynamics and providing more reliable downside risk estimates.
6 Conclusion and Future Work
This study contributes to the advancement of financial volatility modeling by proposing a novel hybrid architecture that unifies traditional econometric structure with modern deep learning techniques. Specifically, we introduce an integrated GARCH-GRU model, which structurally embeds the GARCH(1,1) volatility dynamics directly into the recurrent computations of a GRU cell through a dedicated multiplicative gate. This architectural design enables the model to simultaneously capture key volatility stylized facts, such as clustering and persistence, while also learning complex nonlinear temporal dependencies inherent in financial time series.
Empirical evaluations conducted on major financial indices, S&P 500, Dow Jones Industrial Average, and NASDAQ, demonstrate the superior performance of the proposed framework. The integrated GARCH-GRU model consistently achieves lower forecasting errors (MSE and MAE) compared to a range of baseline models, including standalone neural networks (GRU, LSTM, Transformer), traditional GARCH-family models (GARCH(1,1), GJR-GARCH), and several recently proposed hybrid variants. The model not only offers improved statistical accuracy, but also shows computational efficiency, requiring less than 40% of the training time needed by its LSTM-based counterpart, making it particularly attractive for high-frequency or large-scale financial applications.
In addition to point forecast accuracy, we demonstrate the practical utility of the GARCH-GRU framework in risk management applications through Value-at-Risk (VaR) forecasting. The model produces the lowest violation ratio among all candidates in a one-day-ahead VaR backtesting experiment, aligning closely with the nominal risk level and outperforming both traditional and hybrid benchmarks. This result affirms the model’s ability to deliver reliable tail risk estimates and highlights its relevance for real-world deployment in regulatory or capital allocation settings.
Overall, the proposed integrated GARCH-GRU model demonstrates how a tightly coupled, end-to-end integration of econometric structure and neural network architecture can lead to substantial improvements in both forecasting accuracy and model interpretability. It advances beyond conventional hybrid approaches that treat econometric and machine learning components in isolation, offering a unified framework in which both are optimized jointly. This design enables the model to harness the complementary strengths of both paradigms, combining the statistical rigor of GARCH processes with the nonlinear modeling flexibility of deep learning, thereby yielding a more powerful and principled approach to volatility forecasting.
One current limitation of the proposed framework lies in the computational inefficiency introduced by the use of custom recurrent cell implementations. Although these customized cells have been functionally validated and the GARCH-GRU model demonstrates superior efficiency relative to the GARCH-LSTM, requiring significantly less training time, the reliance on non-native deep learning components incurs runtime overhead compared to standard, highly optimized neural modules. This gap becomes more pronounced in large-scale or low-latency applications. To address this, future work will focus on improving computational efficiency through code-level optimizations, operator fusion techniques, and hardware-aware implementation strategies, enabling the model to scale more effectively while retaining its modeling advantages.
Looking ahead, the flexibility of the proposed framework opens several promising directions for future research. One natural extension involves coupling the GRU architecture with asymmetric GARCH variants, such as EGARCH or GJR-GARCH, to better capture leverage effects and asymmetries in volatility responses to positive and negative shocks. Another avenue lies in the integration of attention mechanisms, which could enhance the model’s ability to focus on relevant historical patterns, particularly in high-frequency financial forecasting where data are noisy and regime-dependent. In addition, the framework could be extended to incorporate regime-switching dynamics, allowing it to model structural breaks and transitions between different volatility regimes over time. These enhancements would further deepen the integration of econometric rigor and machine learning flexibility, resulting in models that are not only highly interpretable and scalable, but also robust to the complexities of modern financial markets.
Appendix A In-Sample Analysis - GARCH Parameter Estimation Results
Table 5: In sample analysis - Parameters Estimation for 1-day ahead Forecast
Model eh_GARCH_GRU eh_GARCH_LSTM bl_GARCH_LSTM pl_GARCH_LSTM GARCH(1,1) GJR-GARCH
S&P 500 $\omega_{0}$ 0.7217 (5.450e-02) 0.6085 (3.490e-02) 0.3525 (1.495e-01) 0.0365 (8.248e-03) 0.0344 (8.969e-03) 0.0342 (6.468e-03)
$\alpha$ 0.3063 (9.290e-02) 0.0636 (1.070e-02) 0.0661 (5.810e-02) 0.1697 (2.513e-02) 0.1597 (2.644e-02) 0.0000 (3.487e-02)
$\beta$ 0.6837 (9.290e-02) 0.9264 (1.070e-02) 0.7610 (1.115e-01) 0.7914 (2.435e-02) 0.8051 (2.709e-02) 0.8270 (2.697e-02)
$\gamma$ 0.2642 (4.517e-02)
DJI $\omega_{0}$ 0.6920 (6.410e-02) 0.6416 (4.930e-02) 0.3485 (1.624e-01) 0.0342 (7.130e-03) 0.0316 (7.373e-03) 0.0317 (5.663e-03)
$\alpha$ 0.6283 (1.903e-01) 0.0505 (9.900e-03) 0.1108 (6.090e-02) 0.1708 (2.534e-02) 0.1656 (2.644e-02) 1.3416e-10 (2.789e-02)
$\beta$ 0.3617 (1.903e-01) 0.9395 (9.900e-03) 0.6821 (9.290e-02) 0.7892 (2.454e-02) 0.7998 (2.629e-02) 0.8275 (2.607e-02)
$\gamma$ 0.2576 (3.799e-02)
NASDAQ $\omega_{0}$ 0.7276 (3.400e-02) 0.5710 (5.560e-02) 0.4508 (1.915e-01) 0.0530 (1.150e-02) 0.0487 (1.229e-02) 0.0521 (9.666e-03)
$\alpha$ 0.6058 (7.090e-02) 0.0752 (3.120e-02) 0.0874 (5.420e-02) 0.1340 (2.075e-02) 0.1252 (2.141e-02) 0.0000 (4.129e-02)
$\beta$ 0.3842 (7.090e-02) 0.9148 (3.120e-02) 0.7615 (8.140e-02) 0.8189 (2.356e-02) 0.8327 (2.566e-02) 0.8372 (3.103e-02)
$\gamma$ 0.2210 (4.238e-02)